OpenStudy (anonymous):

A parking lot charges $3 to park a car for the first hour and $2 per hour after that. If you use more than one parking space, the second and each subsequent car will be charged 75% of what you pay to park just one car. If you park 3 cars for t hours, which function gives the total parking charge?

OpenStudy (anonymous):

A f(t) = 3(3 + 2(t − 1)) B f(t) = (3 + 2t) + 0.75 2(3 + 2t) C f(t) = (3 + 2(t − 1)) + 0.75 2(3 + 2(t − 1)) D f(t) = (3 + 2t) + 0.75(3 + 2t) + 0.75 0.75(3 + 2t) E f(t) = (3 + 2(t − 1)) + 0.75(3 + 2(t − 1)) + 0.75 0.75(3 + 2(t − 1)) i dont understand